# New Factorizations of Yang-Mills Amplitudes

### v1

@@ Discussion
In a companion paper Cachazo and Zhang , we will provide proof of our new factorization formulas using the Cachazo-He-Yuan formalism Cachazo et al. (2014a). Under the constraint of vanishing Mandelstam variables, one can prove that all solutions to the scattering equations Cachazo et al. (2014b, c) become singular, making it evident why the amplitudes for YM theory (and many other theories such as GR, in addition to scalar theories) vanish when certain Lorentz products are set to zero. We will then demonstrate how these singular solutions independently contribute to yield new factorization formulas for YM theories when Lorentz products involving polarizations are turned back on for cases $m=1,2,3,4$.
Returning to our current work, continuing to derive explicit reductions of the YM amplitudes $F_{m,n}(i,j)$ for $m=5$ and higher is valuable, particularly for studying cases with $n\geq 12$. It would be excellent to interpret the ratios of Mandelstam variables $X(s,\rho)$ in (Further Factorizations) as amplitudes or currents for certain theories, or even to identify a Feynman rule for generating them.
Understanding the new factorization at the level of Feynman diagrams is another intriguing avenue. This approach would help elucidate how each summand in our factorization summations operates independently within various Feynman diagrams, thereby providing insights into the physical significance of the hidden zeros in YM amplitudes Rodina (2024).
While this paper focuses on YM amplitudes, it is equally valuable to investigate potential new factorizations for other theories, such as GR, and their extensions in string theories, both bosonic and supersymmetric Green et al. (1988); Mafra et al. (2013); He and Zhang (2017); He et al. (2019). By examining the behavior of their amplitudes around the zeros, we could derive new factorization formulas for these theories. Additionally, exploring their connections to double copy constructions Kawai et al. (1986); Bern et al. (2008) and soft theorems Weinberg (1965) would be highly beneficial. Extending this study from tree-level to loop-level amplitudes is another promising direction. By identifying zeros in loop integrands, we can further explore their new factorization properties. Further insights can be gained from studying the curve integral formalism for ${\rm Tr}(\phi^{3})$ amplitudes, which provides a framework for deriving YM amplitudes at all loop levels via the so-called “scalar scaffolding” procedure Arkani-Hamed et al. (2023a, c, d, b, 2024); Arkani-Hamed and Figueiredo (2024a, b); De et al. (2024).
@@ Acknowledgements.
We especially thank Freddy Cachazo for suggesting this study and for his collaboration on related projects. Our thanks also go to Jin Dong and Song He for their insightful discussions and constructive feedback on our manuscript. The research of Y.Z. was supported in part by a grant from the Gluskin Sheff/Onex Freeman Dyson Chair in Theoretical Physics and by Perimeter Institute. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.

### v3

@@ Discussion
In the companion paper Zhang (2025), we will provide proof of our new factorization formulas using the Cachazo-He-Yuan formalism Cachazo et al. (2014a). Under the constraint of vanishing Mandelstam variables, one can prove that all solutions to the scattering equations Cachazo et al. (2014b, c) become singular, making it evident why the amplitudes for YM theory (and many other theories such as GR, in addition to scalar theories) vanish when certain Lorentz products are set to zero. We will then demonstrate how these singular solutions independently contribute to yield new factorization formulas for YM theories when Lorentz products involving polarizations are turned back on.
Returning to our current work, it would be highly valuable to interpret the ratios of Mandelstam variables $X(s,\rho)$ in (Further Factorizations) as amplitudes or currents associated with specific theories, or even to uncover a Feynman rule that generates them. Exploring the new factorization at the level of Feynman diagrams offers another promising direction. This perspective could shed light on how each summand in the factorization operates independently within different Feynman diagrams, providing deeper insights into the physical significance of the hidden zeros in YM amplitudes Rodina (2025); Zhou (2024); Huang et al. (2025)  [^4] See recent progress at loop level in Backus and Rodina (2025).
While this paper focuses on YM amplitudes, it is equally valuable to investigate potential new factorizations for other theories, such as GR, and their extensions in string theories, both bosonic and supersymmetric Green et al. (1988); Mafra et al. (2013); He and Zhang (2017); He et al. (2019). By examining the behavior of their amplitudes around the zeros, we could derive new factorization formulas for these theories. Additionally, exploring their connections to double copy constructions Kawai et al. (1986); Bern et al. (2008) and soft theorems Weinberg (1965) would be highly beneficial. Extending this study from tree-level to loop-level amplitudes is another promising direction. By identifying zeros in loop integrands, we can further explore their new factorization properties. Further insights can be gained from studying the curve integral formalism for ${\rm Tr}(\phi^{3})$ amplitudes, which provides a framework for deriving YM amplitudes at all loop levels via the so-called “scalar scaffolding” procedure Arkani-Hamed et al. (2024a, 2023b, 2023c, 2023a, b); Arkani-Hamed and Figueiredo (2024a, b); De et al. (2024); Arkani-Hamed et al. (2024c).
@@ Acknowledgements.
We are deeply grateful to Freddy Cachazo for suggesting this study, providing invaluable guidance during its early stages, and offering continued support and encouragement throughout the project. Our thanks also go to Qu Cao, Jin Dong, Song He and Bruno Umbert for their insightful discussions and constructive feedback on our manuscript. The research of Y.Z. was supported in part by a grant from the Gluskin Sheff/Onex Freeman Dyson Chair in Theoretical Physics and by Perimeter Institute. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.